OpenStudy (anonymous):

Rectangular boxes with a volume of 10m^3 are made of two materials. The materials for the top and bottom of the box costs $10/m^2 and the material for the sides of the box costs $1/m^2. What are the dimensions of the box that minimize the cost of the box?

OpenStudy (anonymous):

This is what I have so far: Volume = L*W*H = 10m^3 Cost = 2*W*H + 2*L*H +20*W*L I know I need to minimize the cost function, but I am not how to create a cost function in two variables.

OpenStudy (anonymous):

got it thanks